### v1

# User Centric Semantic Communications

Abstract
Current studies on semantic communications mainly focus on efficiently extracting semantic information to save bandwidth between a transmitter and a user. Although significant process has been made, a fundamental design problem in the semantic communications is that the semantic information is extracted based on certain criteria at the transmitter side along, without considering the user’s actual requirements. As a result, the current semantic communications may cause the loss of critical information that is of primary interest to the user. In such cases, the semantic transmission makes no sense to the user since all the received information is irrelevant. To solve this problem, this paper presents a user centric semantic communication system, where the user sends its request for the desired semantic information to the transmitter at the start of each transmission. Then, the transmitter extracts the required semantic information accordingly. A key challenge is how the transmitter can understand the user’s requests for semantic information and extract the required semantic information in a reasonable and robust manner. We solve this challenge by designing a well-structured framework and leveraging off-the-shelf products, such as GPT-4, along with several specialized tools for detection and estimation. Evaluation results demonstrate the feasibility and effectiveness of the proposed user centric semantic communications.
I Introduction
The concept of semantic communications was originally introduced by Warren Weaver in 1953, focusing on how precisely do the transmitted symbols convey the desired meaning. By concentrating on semantic content, the number of transmitted symbols can be significantly reduced, thereby increasing transmission efficiency in wireless communications. The semantic communications examine the alignment between the transmitter’s intended meaning and the user’s interpretation of the received symbols[1]. A key problem left unsolved in [1] is how to select transmitted symbols that best represent the transmitter’s intended meaning while ensuring alignment with the user’s interpretation.
In [2] and its subsequent works [3, 4, 5, 6, 7], machine learning (ML) techniques such as autoencoders[8] and transformers[9] act as key enablers for the semantic communications. Typically, a neural network-based encoder at the transmitter extracts essential features of the input data, where the dimension of the features is significantly lower than that of the input data. The extracted essential features are defined as the semantic information of the input data [2, 3, 4, 5, 6, 7], which represents the transmitter’s intended meaning. A neural network-based decoder at the user side performs interpretation on the received semantic information by reconstructing the original input data from the transmitter based on it. An end-to-end pre-training is performed beforehand to improve the alignment between the transmitter’s intended meaning and the user’s interpretation.
The current semantic communication framework[2, 3, 4, 5, 6, 7] may lead to the loss of critical information for the user. As illustrated in Fig. 1, although the main content of a video frame is preserved at the user side during the semantic communications, critical details for the user such as a license plate number have been lost. This lost information might be the only detail the user intends to extract from the video frame. As a result, this semantic transmission makes no sense to the user since all the received information is irrelevant, as the information the user cares about has been lost. The loss of critical information problem arises partly due to the out-of-distribution challenge inherent in the ML models when the distribution of an unseen task is different from the distribution of the training data. In this case, the current framework inadvertently discarding subtle yet important information to the user. The out-of-distribution challenge frequently occurs in practice since it is impossible to train an ML model for every possible data distribution.
(a) Original video frame with clear license plate number “BC54950”.
(b) Reconstructed video frame with lost license plate information.
Figure 1: An autoencoder with a compression rate of 0.5 used in semantic communications fails to recover the license plate number of interest to the user due to the out-of-distribution problem. The peak signal-to-noise ratio (PSNR) of (b) is 15. 97 dB. The autoencoder is pre-trained on the MNIST dataset[10] which consists of 70000 handwritten digits images.
The above observation illustrates a fundamental design flaw in the current semantic communication framework. That is, the current framework extracts the so-called semantic information (i.e., compresses input data) based on certain criteria at the transmitter side along, such as PSNR in image transmissions, and does not take the user’s actual requirement into account. If the transmitter were aware of the user’s information needs in advance, it could take actions to prevent the loss of the user’s required information.
In this paper, we present a user centric semantic communication system, where fulfilling the user’s information requirements is central to each semantic transmission. In the proposed user centric system, the semantic information is redefined as the specific information that user intends to get through a particular data transmission. For example, through the transmission of the video frame in Fig. 1, the user wants to get the license plate number. Thus, the license plate number is the semantic information in this video frame at this particular transmission. In user centric semantic communications, each transmission is initialized by the user. Specifically, at the start of each transmission, the user sends its request on the semantic information to the transmitter. Then, the transmitter extracts the related semantic information and sends it to the user.
We restrict our discussion on the semantic communication in scenarios where the original data at the transmitter is surveillance videos, as the user typically have clear requests for semantic information on the surveillance videos, such as checking traffic situations on a particular road. To achieve the user centric communication system, there are at least two significant challenges.
The first significant challenge is how the transmitter can understand the user’s requests for semantic information and extract the required semantic information from the surveillance video accordingly. To address this, we leverage a large language model (LLM) to understand the user’s requests (termed as tasks), and perform task planning. Specifically, we build a system which consists of a LLM agent and several specialized tools such as an object detection tool. Based on its understanding of the semantic information request, the LLM determines which specialized tools to use and the order of their execution. Then, guided by the task planning, the specialized tools work together to extract the semantic information. Finally, the required semantic information is sent back to the user in text form.
The second significant challenge is the possibility that the LLM’s task planing ability is not robust such that the generated task plan is not reasonable sometimes, and thus cannot fulfill the user’s request directly. To make the user still get the required semantic information, a simple but inefficient way is to have the transmitter send the original video to the user, which can lead to significant bandwidth waste. To solve this challenge, we propose a task reflection process where the LLM assesses whether the current task plan can meet the user’s request on semantic information. If the task reflection result is negative, we alternatively ask the LLM and the specialized tools to select the most relevant video frames related to the user’s request from the surveillance video, and send them back to the user. Upon receiving the video frames, the user can get the required semantic information from the video frames by themselves.
We implement the user centric semantic communication system using the off-the-shell LLM product GPT-4 released by OpenAI [11], along with several specialized tools that have specific detection or estimation capabilities. A demo video is available at [12]. We invite 18 postgraduates from 5 Universities to evaluate the overall system’s performance through traffic surveillance videos. The results show that our system can fulfill 83.90% of the users’ requests for semantic information from the traffic surveillance videos. In addition, compared with the traditional system where the whole videos are transmitted for information extraction at the user side, our system can reduce the number of transmitted video frames by 81.70%, and video data size by 66.33%.
In summary, the proposed user centric semantic communication system successfully meets users’ requirements for semantic information in most cases while significantly reducing the transmitted data size. Second, the system is independent of the data distributions of the surveillance video, as the LLM and specialized tools are off-the-shelf products, meaning we do not need to train them using our data. Third, while we discuss the user centric semantic system in the context of surveillance videos, the framework and methods have the potential to be applied to scenarios where users have clear requests for semantic information from the multimodal data stored at the transmitter.
II Overview of Framework
In this section, we overview the framework of the proposed user centric semantic communication system as shown in Fig. 2. The user and transmitter are located separately and communicate with each other through a wireless network. We focus specifically on the semantic information transmission in the context of surveillance video, since the user always has clear requirements on the semantic information regarding the surveillance videos. The communication is triggered by the user. Specifically, at the start of each transmission, the user sends its request on the semantic information of the surveillance video to the transmitter. Then, the transmitter extracts the related information from the surveillance video, and sends the information to the user.
Figure 2: Overall framework of the proposed user centric semantic communication system.
II-A Transmitter Side
We begin by introducing the transmitter components in the framework shown in Fig. 2. The surveillance videos, e.g., traffic surveillance videos, that are concerned by the user are stored at the transmitter side. Apart from this, the transmitter is connected with an LLM, and a toolbox consisted with several specialized tools. In this paper, GPT-4, released by OpenAI [11], serves as the LLM. In addition, denote the toolbox by
[…]
where $T_{n}$ denotes a specialized tool, $n=1,\dots,N$, and $N$ is the number of tools in the toolbox. Each specialized tool is responsible for one specific detection or estimation task during the semantic information extraction. In Section V, we evaluate the proposed system through traffic surveillance videos, where we apply 8 specialized AI tools to help extract semantic information. The toolbox is shown in Table I.
TABLE I: Tools in the Toolbox
Index Tool Name Brief Description
1 Object Detection[13] Detects 80 different objects, e.g., person, bicycle, hydrant, and motorcycle.
2 Vehicle Detection[14] Detects 9 types of vehicles, e.g., sedan, SUV, bus, pickup, and truck.
3 License Plate Detection[14] Detects license plates on vehicles.
4 Traffic Sign Detection[15] Detects 45 types of traffic signs, e.g, bicycle path, and pedestrian crossing.
5 Vehicle Motion Detection Detects if vehicles are moving or not in a video.
6 Lane Number Detection[16] Estimates the number of lanes in a road.
7 Traffic Flow Estimation Estimates the number of different vehicles within a time interval in a video.
8 Vehicle Density Estimation Estimates the vehicle density in a specific image.
The user’s request could be complex and diverse, making it impossible to fulfill with a single specialized tool. The transmitter should first understand the user’s request, and then breaks down the complex request into several smaller and manageable steps (referred to as task planning ), in which each step is executable by a specialized tool. Executing these steps sequentially ensures effective task completion. The transmitter achieves this by leveraging LLM. Specifically, the transmitter relays the user’s request to the LLM, prompting it to perform task planning for semantic information extraction. A key challenge in task planning is ensuring that the LLM can perform planning in a reasonable and robust manner, so the resulting task plan reliably fulfills the user’s request. We will show the task planning design in Section III-A.
Then, the transmitter performs task reflection to determine whether we use the current task plan to extract semantic information for the user, or not. In the case the transmitter uses a wrong task plan, the user will not get the required semantic information. In this case, alternative actions should be taken promptly, preventing any negative impact on the system’s performance. Therefore, the task reflection plays a crucial role in our system. The design of the task reflection is shown in Section III-B. There are two outcomes from the task reflection:
1) The current task plan can fulfill the user’s request. In this case, task execution is performed step-by-step, generating the required semantic information in text form. The transmitter then sends this text-based information to the user. Please refer to Example 1.
2) The current task plan cannot fulfill the user’s request. In this case, another round of task planning is triggered. In the case all the task plans cannot fulfill the user’s request, to make the user still get the required semantic information, a simple but inefficient way is to have the transmitter send the original video to the user, which can lead to significant bandwidth waste. To address this issue, we propose leveraging the LLM and specialized tools for frame selection, i.e., select the most relevant video frames related to the user’s request from the surveillance video. Through the selected frames, the user can get the required information by themselves. We will show how to achieve this in Section IV. Last, the transmitter sends the related video frames to the user. Please refer to Example 2.
The overall processes at the transmitter are summarized in Algorithm 1, where $\bar{N}$ is a parameter that controls the maximum number of iterations.
Algorithm 1 Overall Process at the Transmitter
1:Input: $n=1$, $T$, and iteration control parameter $\bar{N}$.
2:while $n\leq\bar{N}$ do
3:     Perform task planning and task reflection.
4:     if Task plan can fulfill request then
5:         Perform task execution.
6:         Send the text-based information to the user.
7:         Terminate the Algorithm.
8:     else
9:         $n\leftarrow n+1$
10:     end if
11:end $\triangleright$ Fail to find a task plan
12:Perform frame selection and send the frames to user.
Example 1 (Fulfillment). User’s request: “Is there a traffic jam in the video”? Transmitter’s feedback : “I can fulfill your request on the semantic information directly. Based on the analysis of the video, there is no traffic jam detected. The traffic density remains 0.0 for most of the video. Towards the end, it slightly increases to 0.01, but this is still considered light traffic. Therefore, we can confidently say that there is no traffic jam in the video”.
Example 2 (Non-fulfillment). User’s request: “Did an accident happen in the video”? Transmitter’s feedback : “I cannot fulfill your request on the semantic information directly. However, I send the most relevant video frames to you. You may get the required semantic information from the attached video frames (as shown in Fig. 3)”.
II-B User Side
The user’s role is to send requests for semantic information about the surveillance videos of interest and to interpret the feedback from the transmitter: If the feedback from the transmitter is in text form, the user can obtain the required semantic information directly; If the feedback is in image form, the user can get the required semantic information by interpreting the images. For example, in the context of Example 2, by interpreting the video frames from the Fig. 3, the user can determine that an accident has occurred from the second selected video frame shown in Fig. 3(b).
(a) Video Frame 1.
(b) Video Frame 2.
Figure 3: Selected video frames where the objects highlighted in red boxes may be of interest to the user.
III Task Planning and Reflection
III-A Task Planning
We first introduce the design in task planning through LLM at the transmitter. Task planning with LLMs involves breaking down complex user requests into smaller and manageable steps. However, relying solely on the LLM to perform task planning without any guidance or structured design is insufficient. Specifically, simply inputting the user’s request into the LLM and expecting it to generate a reasonable plan often fails because LLM has a limited understanding of the user’s specific requirements, which can result in impractical or unfeasible steps within the task plan. Additionally, LLM struggles with maintaining long-term dependencies and complex sequencing of steps, resulting in disjointed or out-of-order actions[17]. For example, the first step of the task plan might be mistakenly placed as the third step, rendering the entire plan unreasonable.
To solve the above problems, we have two particular designs for task planning. First, we fix the overall task planning to three key steps:
[…]
The description of each step is as follows:
  • Video Sampler: LLM samples the input video frames that are used in the next step, according to the user’s request and video frame information (e.g., frame rate and total number of frames). For example, the user may be interested in the semantic information on a particular time of the video.
  • Tool Selection: LLM selects only one appropriate tool in the toolbox $T$ in (1), according to the user’s request. Then, the tool is executed to perform detection or estimation, according to the understanding of the LLM.
  • Analysis: LLM responds to the user’s request based on the execution results from Video Sampler and Tool Selection
In summary, to ensure the task plan is coherent across steps, we apply the LLM only to perform roles in each step, as the overall structure of the plan has been fixed in advance.
Second, a key step in the above task planning design is enabling the LLM to select one suitable specialized tool from the toolbox $T$ during the Tool Selection step. To facilitate this, we first provide the LLM with the toolbox $T$, allowing it to choose from existing tools rather than generating one. Additionally, to ensure the LLM selects one appropriate tool, we include a detailed description for each tool in the toolbox $T$, which encompasses: 1) the functions the tool can perform; 2) the functions it cannot perform; 3) the output format, along with an example. The detailed description of each tool is shown in Appendix -C. Below is an example of the tool Vehicle Density Estimation.
Example 3 (Detailed Description of Vehicle Density Estimation).
1. The algorithm can a) compute vehicle density in a particular video frame; b) indicate traffic states in a video frame. For example, higher vehicle density indicates traffic congestion or heavy traffic flow on the roads. 2. The algorithm cannot perform detection on a specified location in the video. For example, at the crossroad, or along the road. 3. The output is the vehicle density, represented as a number between 0 and 1, with higher values indicating heavier traffic. For example, the output is: 0.23.
We implement the above task planning through the prompt engineering technique[18] in LLM, which is summarized in Algorithm 2.111When implementing Line 2 of Algorithm 2, we included a detailed description of the three steps involved in executing the prompt. To save space, we omit them here. Through the prompt, along with the other processes in the framework shown in Fig. 2, the Algorithm 2 can output a task plan for the user’s request. For example, the task plan for the request “Is there a traffic jam in the video?” is shown in Line 4 in Algorithm 2.
Algorithm 2 Prompt in Task Planning
1:Input: {User’s Request}, and {Toolbox $T$ with Detailed Description}.
2:You are equipped with a toolbox $T$. To fulfill the user’s request, you need to plan first. The plan consists of following three steps in order: Video Sampler, Tool Selection, and Analysis.
3:In the Tool Selection step, you must select one proper tool from toolbox $T$ that can help to perform detection or estimation in the video. You should only use one tool from the toolbox each time in this step.
4:Your response on Task Planning should be logical and step by step, with each step divided by the symbol $\mid$”. Here’s an example: “Video Sampler $\mid$ Vehicle Density Estimation $\mid$ Analysis”.
Algorithm 3 Prompt in Task Reflection
1:Input: {User’s Request}, {Current Task Plan}, and {Toolbox $T$ with Detailed Description}.
2:The overall framework to solve this request consists of three steps: Video Sampler, Tool Selection, and Analysis.
3:Do you think the current task plan {Current Task Plan} can exactly address the user’s request? You must answer with “Yes” or “No”. Also, you must provide your explanations on this answer.
III-B Task Reflection
After task planning, the transmitter should reflect whether the current task plan can fulfill the user’s request. This step is crucial for the overall system because if the transmitter determines that the current plan cannot satisfy the user’s needs, alternative actions can be taken promptly, preventing any negative impact on the system’s performance. There are two scenarios that may lead to the user’s request not being fulfilled. First, the LLM might select the wrong tool during the Tool Selection step. Second, the toolbox may lack sufficient tools, forcing the LLM to choose an inappropriate tool from the available options. This is a practical concern, given the diverse nature of user’s requests.
The transmitter performs the task reflection by asking the LLM to reflect the current task plan. The prompt that guides the LLM to perform task reflection is summarized in Algorithm 3. Below are two examples.
Example 4 (Positive Reflection Result). User’s request: “Is there a traffic jam in the video”? Current task plan : “Video Sampler $\mid$ Vehicle Density Estimation $\mid$ Analysis”. Task reflection: Yes, this task plan can address the user’s request. The Video Sampler step will help to identify the relevant video frames which is crucial in narrowing down the scope of the analysis to the relevant parts of the video. The Tool Selection step, using the Vehicle Density Estimation tool, will extract the necessary information from the selected video frames. This algorithm is particularly suited to the task as it can compute the vehicle density in a particular video frame, which is a direct indicator of traffic congestion. Finally, the Analysis step will summarize the results from the Vehicle Density Estimation, providing a clear and concise output for the user.
Example 5 (Negative Reflection Result). User’s request: “Did an accident happen in the video”? Current task plan : “Video Sampler $\mid$ Vehicle Motion Detection $\mid$ Analysis”. Task reflection: No, the task plan cannot exactly address the user’s request. The user’s request is to determine if an accident happened in the video. While the framework can select frames, detect vehicle motion, and analyze the results, it does not have the capability to specifically identify an accident.
Last, according to Algorithm 1, if the task reflection result is negative, the transmitter should generate an alternative task plan. The task re-planning process is the same as the task planning process introduced in (2), except that during the Tool Selection step, the LLM selects a different appropriate tool from the remaining options in the toolbox $T$. Tools that have already been selected are removed from the toolbox during this step. In addition, in Algorithm 1, the maximum value of the iteration control parameter $\bar{N}$ can be set to $N$ since each iteration in Algorithm 1 essentially selects a different tool from the toolbox $T$, which contains a total of $N$ tools.
Algorithm 4 Prompt in Frame Selection
1:Input: {User’s Request}, and {Execution Results of Selected Tool}.
2:The {Execution Results of Selected Tool} only provides partial information and cannot fully fulfill the user’s request. Based on {Execution Results of Selected Tool}, you need to select the most relevant frames that may help to fulfill the user’s request.
3:Do not answer user with this tool’s execution result. You should only output the identification numbers (IDs) of the relevant frames.
4:For example, the user wants to know the number of motorcyclists wearing helmet in the whole video. However, the selected tool can only detect the motorcyclists appears in the video. In this case, you should output the frame IDs of frames with motorcyclists.
IV Frame Selection
When none of the task plans can satisfy the user’s request, according to Algorithm 1, the transmitter should select the most relevant video frames from the surveillance video. This frame selection process can be accomplished by leveraging the LLM and specialized tools.
IV-A Tool Selection
Consider the following two representative user’s requests that none of the task plans can satisfy by applying the toolbox in Table I. The first request is: “Did an accident happen in the video?” We cannot satisfy this request as we lack a tool to detect accident; The second request is “How many motorcyclists wearing helmet in the whole video?” Again, we cannot satisfy this request as we lack a tool to detect motorcyclists while wearing helmets, although we do have a tool to detect motorcyclist. In summary, a reason all task plans fail to meet user’s requests is that the tools in the toolbox are not enough. This is a practical case since the user’s request is diverse while the tools are always limited.
Our strategy is, we ask the LLM to extract one key term from the user’s request, and map it to a tool in the toolbox $T$. In this case, this tool can assist us by selecting relevant video frames. To illustrate this, for the request “Did an accident happen in the video?”, if the key term of this request could be “accident”, then we can apply Vehicle Motion Detection to locate the video frames with non-moving vehicles, since there is a high probability that the vehicles involved in an accident are not moving. As shown in Example 2 and Fig. 3, through the two selected video frames with non-moving cars, the user can determine there is an accident in Fig. 3(b); For the request “How many motorcyclists wearing helmet in the whole video?”, if the key term of this request could be “motorcyclist”, then we can apply Object Detection to locate the video frames with motorcyclists. The user can count the number of motorcyclists wearing helmets by themselves through the selected video frames.
[…]
Example 6. Tool Name: Vehicle Motion Detection. Labels: Accident, Collision; Tool Name: Vehicle Detection. Labels : sedan, SUV, van, hatchback, MPV, pickup, bus, truck, and estate.
IV-B Overall Process
The overall process of the frame selection is as follows:
[…]
The Video Sampler is same as that shown in III-A. For the Frame Selection, we input the execution results from the selected tool to the LLM, and ask the LLM to select the frames that may help to fulfill the user’s request. The prompt that guides the LLM to perform frame selection is summarized in Algorithm 4.
V System Performance
V-A Implementation
We implement the user-centric semantic communication system in the context of traffic surveillance videos, using a computer equipped with an NVIDIA RTX 4070 GPU, and a cell phone. First, the computer acts as the transmitter. It connects remotely to OpenAI’s GPT-4 and stores the surveillance videos locally. Additionally, the 8 specialized AI tools in Table I are installed on the computer and utilize its local resources for detection or estimation tasks. Second, both the cell phone and the computer have WeChat installed (an instant messaging application). The user sends their semantic information requests through WeChat on the cell phone, which are received by the computer. Likewise, the computer sends its feedback—text or video frames—back to the user via WeChat.
A demo video of the system is available at [12].
V-B Evaluation Methodology and Results
Evaluation Datasets. Our dataset consists of 100 traffic surveillance video clips, each lasting roughly 15 seconds. Among them, 47 video clips are sourced from the datasets CityFlow[19] and CityFlow V2[20], and the other 53 video clips are made by us. In addition, our dataset also consists of 700 possible semantic information requests on the traffic surveillance videos, with 300 labeled “Yes” and 400 labeled “No”. This means, the ground truth is that 300 user’s requests can be fulfilled through our system while 400 cannot be satisfied by any of the generated task plans in our system. We denote Req Y for a request labeled “Yes”; Req N for a request labeled “No”; and Req YN for a general request that is either Req Y or Req N. The dataset has been released to[21]. Last, each video clip is associated with 7 semantic information requests in average.
Figure 4: Performance of Task Reflection, where $\bar{N}$ in x-axis is a parameter that controls maximum number of iterations in Algorithm 1.
Performance of Task Reflection in Algorithm 1. We test its performance through the 700 requests in our dataset. Specifically, we input the 700 requests into Algorithm 1, generating corresponding task plans. Then, the task reflection assesses the feasibility of these plans. An accurate evaluation of Req Y occurs when the task reflection believes Req Y can be fulfilled by the task plan; if otherwise, the evaluation is inaccurate. An accurate evaluation of Req N occurs when the task reflection consistently believes Req N cannot be fulfilled by the task plans generated during the $\bar{N}$ iterations in Algorithm 1; if otherwise, the evaluation is inaccurate. The accurate ratio of Req Y is thus defined as the ratio of accurate evaluations of Req Y to the total number of Req Y; the accurate ratio of Req N is defined as the ratio of accurate evaluations of Req N to the total number of Req N; and the accurate ratio of Req YN is defined as the ratio of accurate evaluations of both Req Y and Req N, to the total number of all requests.
The evaluation results are shown in Fig. 4. The results show that, when $\bar{N}=3$, the accurate ratio of the three types of requests exceeds 90%. Specifically, for Req Y, the accurate ratio reaches its maximum value when $\bar{N}\geq 3$. This is because trying more tools can indeed increase the task reflection accurate ratio. However, since once the LLM confirms the selection of a tool, the iterations in Algorithm 1 are terminated. Therefore, increasing of $\bar{N}$ beyond 3 does not further improve accurate ratio; For Req N, the results show that increasing $\bar{N}$ will decrease its accurate ratio. This is due to the fact that, when the task reflection process decides a request cannot be fulfilled, it need to check all $\bar{N}$ tools. If it makes a mistake with any of the $\bar{N}$ tools, the task reflection process may incorrectly believe that the request can be fulfilled. Therefore, a larger $\bar{N}$ can increase the probability of such mistakes. Considering the accurate ratio of Req Y and Req N, as a result, the overall accurate ratio Req NY reaches its maximum at $\bar{N}=3$.
Performance of Overall System. We evaluate the overall user centric semantic communication system through the 100 video clips and the associated 700 user’s requests when $\bar{N}=3$ in Algorithm 1. In addition, we invite 18 postgraduates from 5 Universities to test the performance of the overall system. Specifically, for each user’s request, they receive a feedback from the transmitter (in text form or video frame form). Then, we ask them to evaluate if the transmitter’s feedback can fulfill the user’s request. Evaluations are recorded as either “Yes” or “No”. The user success rate in obtaining semantic information is defined as the ratio of “Yes” evaluations to the total number of requests.
We have conducted 3 independent experiments, in which we test the system’s performance over the dataset repeatedly for 3 times. In addition, for each experiment, we invite 3 different groups of postgraduates to evaluate the system’s performance. The average results are shown in Table II. The results show that our system can fulfill 83.90% of the users’ requests overall, with the success rate for Req Y being higher than that for Req N. This is due to the fact that the performance of Frame Selection is not that good. We leave it as a future work to further improve the performance.
TABLE II: User Success Rate in Obtaining Semantic Information
Exp 1 Exp 2 Exp 3 Average
Req Y 90.89% 92.44% 93.56% 92.30%
Req N 76.58% 77.25% 77.41% 77.08%
Req YN 82.71% 85.00% 84.00% 83.90%
TABLE III: Resource Reduction Ratio Compared with Traditional System
Exp 1 Exp 2 Exp 3 Average
Frame Count Reduction Ratio 81.49% 81.37% 82.24% 81.70%
Data Size Reduction Ratio 66.42% 65.52% 67.07% 66.33%
We also compute the resource savings achieved by our system in terms of frame count (i.e., the number of video frames) and data size, compared with the traditional system in which the whole videos are transmitted to the user. A significant reduction in frame count and data size can substantially lower the transmission bandwidth required for wireless communications. In addition, for the transmission of the video frames, we do not apply any compression techniques for further processing.
From Table III, the frame count is reduced significantly. For the data size, although it is reduced by around 66.33%, the reduction ratio is lower than that of the frame count. A main reason is, although a video clip consists of many video frames where each video frame is with a large data size, advanced compression techniques significantly reduce the video’s overall size. Consequently, the compressed video is much smaller than the combined size of all its video frames. However, when individual video frames are extracted, they retain their large data sizes, resulting in a substantial data size. Thus, the date size reduction ratio is lower than the frame count ratio. We plan to address this problem in a further work.
-C Tools and Their Detailed Descriptions
The tools and their detailed descriptions are listed as follows:
  • Object Detection: 1. The algorithm can a) detect 80 different objects including: Person, Bicycle, Car, Motorcycle, Airplane, Bus, Train, Truck, Boat, Traffic Light, Fire Hydrant, Stop Sign, Parking Meter, Bench, Bird, Cat, Dog, Horse, Sheep, Cow, Elephant, Bear, Zebra, Giraffe, Backpack, Umbrella, Handbag, , Suitcase, Frisbee, Skis, Snowboard, Sports Ball, Kite, Baseball Bat, Baseball Glove, Skateboard, Surfboard, Racket, Bottle, Wine Glass, Cup, Fork, Knife, Spoon, Bowl, Banana, Apple, Sandwich, Orange, Broccoli, Carrot, Dog, Pizza, Donut, Cake, Chair, Couch, Potted Plant, Bed, Dining Table, Toilet, TV, Laptop, Mouse, Remote, Keyboard, Phone, Microwave, Oven, Toaster, Sink, Refrigerator, Book, Clock, Vase, Scissors, Teddy Bear, Hair Drier, Toothbrush; b) provide only static analysis of frames in the video, meaning it cannot track the movement of objects. 2. The algorithm cannot a) do detection on a specified location in the video. For example, at the crossroad, or along the road; b) do detection in a specified event in the video. For example, after an accident, after a collision. 3. The output includes the name of the detected object in each frame and the corresponding probability. Here is an example of the output: [(0.98, ‘car’), (0.95, ‘traffic light’), (0.89, ‘traffic light’), (0.87, ‘car’), (0.83, ‘fire hydrant’)].
  • Vehicle Detection: 1. The algorithm can a) detect types of vehicles including sedan, SUV, van, hatchback, MPV, pickup, bus, truck, and estate; b) identify colors of vehicles including yellow, orange, green, gray, red, blue, white, golden, brown, and black; c) provide only static analysis of frames in the video. 2. The algorithm cannot a) track the movement of vehicles or events that occur during the movement of the vehicles; b) do detection on a specified location in the video. For example, at the crossroad, or along the road; c) do detection in a specified event in the video. For example, after an accident, after a collision. 3. The output includes color and type of vehicles in frames of the video. Here is an example of the output: 1: [(‘Color: blue’, ‘Type: sedan’), (‘Color: blue’, ‘Type: hatchback’)], 2: [(‘Color: blue’, ‘Type: sedan’)].
  • License Plate Detection: 1. The algorithm can detect vehicle license plates in videos. This algorithm outputs the detected license plate numbers for each frame of the video. 2. The output includes the license plate numbers and characters. Here is an example of the output: 1: [‘C’, ‘J’, ‘X’, ‘S’, ‘G’].
  • Traffic Sign Detection: 1. This algorithm can detect 45 types of different traffic signs in the video, including: Speed Limit 80, No Bicycles, No U-Turn, Maximum Weight 55t, Speed Limit 60, Pedestrian Crossing, No Honking, Non-motor Vehicle Lane, No Left Turn, Yield, Minimum Speed Limit 80, Height Limit 4m, Motor Vehicle Lane, Speed Limit 70, No Entry, Height Limit 4.5m, No Motorcycles, No Large Buses, No Rickshaws, Crossroad Motor Vehicle Lane, Speed Limit 30, No Motor Vehicles, No Parking (Except for Loading or Unloading), Children Crossing, No Trucks, No Two Specific Vehicles, End of Speed Limit, Speed Limit 20, Maximum Weight 30t, Speed Limit 40, Non-motor Vehicle Lane, Speed Limit 120, Road Work Ahead, Height Limit 5m, Minimum Speed Limit 60, Pedestrians Crossing, Speed Limit 100, Merge Ahead, Minimum Speed Limit 100, No Right Turn, Maximum Weight 20t, Keep Right, No Hazardous Materials, Speed Limit 50. 2. The algorithm cannot a) detect other types of traffic signs; b) detect the location of traffic signs. 3. The output includes the name of the detected traffic sign in each frame and its corresponding probability. Here is an example of the output: (0.95, ‘Speed Limit 70’), (0.86, ‘No Trucks’).
  • Vehicle Motion Detection: 1. The algorithm can detect whether vehicles in the video are moving or not. For a particular vehicle, its motion is determined as follows. First, for the same vehicle, the algorithm detects and computes the occupied areas on two different selected image frames, respectively. Then, the algorithm computes the ratio of the occupied areas. Last, if the ratio is larger than a predefined threshold, the algorithm claims the vehicle is moving; otherwise, the vehicle is not moving. 2. The algorithm cannot a) do detection on a specified location in the video. For example, at the crossroad, or along the road; b) do detection in a specified event in the video. For example, after an accident, after a collision. 3. The output includes two motion states of different vehicles in the video: moving or not moving. Here is an example of the output: [not moving, moving, moving, not moving, moving].
  • Lane Number Detection: 1. The algorithm can detect the number of road lanes of the selected image frames in the video. 2. The algorithm cannot recognize different types of lanes, such as solid lanes, dashed lanes and arrows. 3. The output includes the number of lanes. Here is an example of the output: number of lanes : [3].
  • Traffic Flow Estimation: 1. The algorithm can count the total number of different vehicles over a particular time period in a video. 2. The algorithm cannot a) recognize the path or motion of vehicles; b) do detection on a specified location in the video. For example, at the crossroad, or along the road. 3. The output is the total number of the different vehicles. Here is an example of the output: Total vehicle number: 8.
  • Vehicle Density Estimation: 1. The algorithm can a) compute vehicle density in a particular image frame. It is the ratio of the area occupied by vehicles to the area of the road in an image frame; b) indicate traffic states in a frame. For example, higher vehicle density indicates traffic congestion or heavy traffic flow on the roads. 2. The algorithm cannot perform detection on a specified location in the video. For example, at the crossroad, or along the road. 3. The output is the vehicle density, represented as a number between 0 and 1, with higher values indicating heavier traffic. For example, the output is: 0.23.
-D Tool and Their Labels
The tools and the corresponding labels used in Frame Selection in Section IV are listed as follows:
  • Object Detection: Person, Bicycle, Car, Motorcycle, Airplane, Bus, Train, Truck, Boat, Traffic Light, Fire Hydrant, Stop Sign, Parking Meter, Bench, Bird, Cat, Dog, Horse, Sheep, Cow, Elephant, Bear, Zebra, Giraffe, Backpack, Umbrella, Handbag, , Suitcase, Frisbee, Skis, Snowboard, Sports Ball, Kite, Baseball Bat, Baseball Glove, Skateboard, Surfboard, Racket, Bottle, Wine Glass, Cup, Fork, Knife, Spoon, Bowl, Banana, Apple, Sandwich, Orange, Broccoli, Carrot, Dog, Pizza, Donut, Cake, Chair, Couch, Potted Plant, Bed, Dining Table, Toilet, TV, Laptop, Mouse, Remote, Keyboard, Phone, Microwave, Oven, Toaster, Sink, Refrigerator, Book, Clock, Vase, Scissors, Teddy Bear, Hair Drier, and Toothbrush.
  • Vehicle Detection: Sedan, SUV, van, hatchback, MPV, pickup, bus, truck, and estate.
  • Traffic Sign Detection: Traffic Sign.
  • Vehicle Motion Detection: Accident, and Collision.